# Neutron Reflectometry by Gradient Descent

### v1

@@ Abstract
Neutron reflectometry (NR) is a powerful technique to probe surfaces and interfaces. NR is inherently an indirect measurement technique, access to the physical quantities of interest (layer thickness, scattering length density, roughness), necessitate the solution of an inverse modelling problem, that is inefficient for large amounts of data or complex multiplayer structures (e.g. lithium batteries / electrodes). Recently, surrogate machine learning models have been proposed as an alternative to existing optimisation routines. Although such approaches have been successful, physical intuition is lost when replacing governing equations with fast neural networks. Instead, we propose a novel and efficient approach; to optimise reflectivity data analysis by performing gradient descent on the forward reflection model itself. Herein, automatic differentiation techniques are used to evaluate exact gradients of the error function with respect to the parameters of interest. Access to these quantities enables users of neutron reflectometry to harness a host of powerful modern optimisation and inference techniques that remain thus far unexploited in the context of neutron reflectometry. This paper presents two benchmark case studies; demonstrating state-of-the-art performance on a thick oxide quartz film, and robust co-fitting performance in the high complexity regime of organic LED multilayer devices. Additionally, we provide an open-source library of differentiable reflectometry kernels in the python programming language so that gradient based approaches can readily be applied to other NR datasets.
Keywords Neutron Reflectometry  $\cdot$ Automatic differentiation
@@ 1 Introduction
Neutron reflectometry (NR) is a widely used technique for the nanometer scale characterisation of soft matter [1], thin films [2] and bio-materials [3]. Increasingly, this has involved the study of more complex layered structures and ones that are evolving in time (dynamic). The application of NR has lead to many high-profile breakthroughs in the physical and biological sciences. Polarized neutron reflectivity (PNR) was crucial for understanding the oscillatory exchange coupling behaviour so fundamental to giant magnetoresistance (GMR) [4], a pivotal modern technology. Ongoing work has investigated how surfactant mixtures perform at varying temperatures using surface tension and neutron reflectivity [5], this has been crucial for developing more efficient, low-temperature detergents, vital for reducing the carbon footprint of washing laundry. Ongoing NR studies have also been important for the study of biological membranes [6], their interaction with drugs [7], such as antimicrobial peptides and other antibiotics.
NR is a scattering technique that illuminates a sample of interest using thermal neutrons produced and moderated from either a reactor or a spallation neutron source. NR measures the specular reflectivity $R$ as a function of the momentum transfer $Q$ from counting neutrons using a detector [8]. Theoretically, the forward reflectivity model of complex materials is well understood. Given the physical and apparatus parameters $\theta$, one can appeal to Fresnel reflectivity to compute the predicted specular reflectivity $\hat{R}$. NR is however, inherently an indirect technique, which requires the reflectivity data to be modelled, as it is the absolute square of Fourier transform of the scattering length density, which gives rise to a loss in phase information. The parameters of interest cannot be directly accessed from measurements of the reflectivity. Fitting a parametric model of the sample is the inverse problem at the core of NR measurements and data analysis.
In the four decades since the widespread adoption of NR as a measurement technique, many methods have been proposed to solve the inverse modelling problem. Until recently, advances in NR analysis and fitting have kept close pace with the progress in machine learning (ML). The rapid rise of deep learning, automatic differentiation and specialised computation libraries has revolutionised the scalability of ML in many other fields. Yet, these analysis routes have so far been very slow to translate to NR fitting. Among the leading software packages for NR analysis [9, 10], the underlying fitting routines are often based on traditional data optimisation techniques such as Levenberg-Marquardt (LM) [11, 12], Broyden-Fletcher-Goldfarb-Shanno (BFGS) [13], or heuristic techniques such as differential evolution [14] or particle swarms [15]. Other approaches follow Bayesian methodologies such as Markov-chain Monte-Carlo (MCMC) [16, 17] or nested sampling [18]. The authors are not currently aware of any widely-used NR software that fully utilises the power of modern automatic-differentiation and gradient-based optimisation techniques to directly solve the inverse problem.
Although no direct applications of automatic-differentiation are to be found in the literature, surrogate models have very recently been proposed to replace the inverse problem [19, 20, 21, 22] with ML. These ML models supplant the forward reflectivity model and are trained to predict the parameters of interest directly from the measured $R$. Although such approaches are able to leverage advanced ML technologies, it is inevitable that physical intuition is lost when replacing governing equations with fast machine learning.
Instead, we propose a more robust approach to efficiently solve the NR inverse problem by evaluating the gradient of the forward physical model itself with respect to the parameters of interest. In this paper, we demonstrate that modern tensor computation libraries supporting automatic differentiation are able to quickly compute these derivatives, unlocking the potential to leverage a host of powerful optimisation and inference techniques that remain thus far unexploited in the context of neutron reflectometry. The principal contributions of this work are:
  • The gradient of the forward specular reflectivity of a thin multilayer film (w.r.t the unknown parameters of interest) is computed using automatic differentiation techniques.
  • The gradient information is supplied to cutting edge ML optimisation (stochastic gradient descent) and inference (Hamiltonian Monte-Carlo, variational inference) techniques, previously unexploited in the context of NR inverse problems.
  • The proposed approach is demonstrated in two NR case study examples: A benchmark Quartz film and a co-fit to four LED devices.
@@
  • An open-source software library is provided for fast computation of forward reflectivity and gradients in the python language at github.com/MDCHAMP/refjax.
1.1 NR inverse problem
@@
At the very core of NR data analysis, is the inverse problem of obtaining estimates of unknown parameters of interest pertaining to both the slab model of the sample $\theta_{\text{slab}}$ (layer thicknesses, scattering length densities and roughnesses) and the instrument $\theta_{\text{NR}}$ (scaling/alignment parameters and the background detection level). For notational simplicity, in the description that follows these unknown parameters shall be referred to jointly as $\theta=\{\theta_{\text{NR}},\theta_{\text{slab}}\}$. The Abeles matrix formalism [23] provides a physical model of the reflectivity in terms of these parameters and the momentum transfer $\hat{R}(Q,\theta)$. Predictions from the forward model can be compared to the observed reflectivity in order to determine the values of the unknown parameters. A key challenge in NR, is the non-uniqueness of the inverse problem because the phase information is lost when measuring the reflected neutron intensity. For a particular measured neutron reflectivity measurement $R(Q)$, there may be many possible slab model solutions and therefore many sets of parameters that describe the data equally well. In practice, this is circumvented by initialising optimisation routines in the vicinity of optima that are physically plausible given prior knowledge of the sample under investigation. Such local optimisation is an area in which gradient-based methods excel, particularly in high-dimensions.
@@
In both the optimisation and inference setting, powerful techniques based on automatic differentiation exist in the ML literature for solving inverse problems. Access to the gradients of the error function (from this work) means that these approaches can be readily applied to NR. The remainder of this section will be dedicated to introducing approaches from both categories that are applied in the case studies outlined in this paper.
@@ 1.2 Optimisation for NR by gradient descent
Although several approaches that use approximate gradient information [11, 12, 13] are common in the NR literature, the authors are not aware of any method that explicitly harnesses the analytical gradient with respect to the parameters of interest using automatic differentiation techniques. Methods in this class have seen intensive development since the rise of deep neural networks, particularly stochastic gradient descent methods. Access to gradient information as in this paper will make the application of these schemes very straightforward to the NR community. Efficient implementations of many popular algorithms are readily available in a number of programming languages. The most popular methods (e.g. Adam [24], RMSPROP [25]) make use of gradient averages between iterations and momentum terms that help to overcome local minima. Such approaches help to alleviate some of the inherent shortcomings of these popular gradient descent approaches when applied to NR data analysis, such as becoming trapped in local minima and computational efficiency in the case of high dimensions of very many parameters.
@@ 1.3 Inference for NR by Hamiltonian Monte-Carlo
In the interest of keeping this paper self-contained and in order to appeal to practitioners of NR who may not be familiar with gradient-based sampling and approximate inference schemes, brief overviews of Hamiltonian Monte-Carlo (HMC) and variational inference (VI) are included here. The interested reader is directed to [26] (HMC) and [27] (VI) for more comprehensive tutorial overviews.
@@
Where the differentiation has removed the unknown normalisation constant in the kinetic energy and the ODE now depends on the hyperparameters, likelihood and prior only. As with any other nonlinear ODE it is necessary to integrate forward in time numerically in order to obtain trajectories (and therefore samples from the target posterior distribution). Symplectic integrators are a class of integrator specifically designed to simulate Hamiltonian dynamics, by attempting to conserve (or bound the drift in) the systems total energy through the Hamiltonian. A popular solver in the context of HMC is the leapfrog algorithm [28]. For a given integration period $\Delta t$, the dynamics are updated sequentially as,
@@
Thus, the overall sampling scheme consists of a Markov chain with the following kernel. In each iteration, an initial momentum is drawn, and the integration is run for a number of steps $L$ with a step size of $\Delta t$. Although the position at the end of the integration window should theoretically be a valid sample from the target distribution, in practice, numerical errors in the integration can lead to bias. To overcome this, a Metropolis-Hastings step is typically taken after integration. At the end of the integration window, the last value of the parameter is accepted as the next state of the Markov Chain with probability $\alpha$. It can be shown (in [26] for example) that the correct choice of $\alpha$ to ensure detailed balance in the Markov Chain is given by,
@@
In practice, the performance of the HMC scheme comes to depend strongly on the choice of the hyperparameters $L,\Delta t$ and $M$. However, the adaptive specification of these parameters is a well studied problem. Variants of the standard HMC algorithm such as the no U-turn sampler (NUTS) [29] have shown incredible robustness to a wide range of posterior geometries. NUTS is now a standard approach in Bayesian inference and implementations can be found in many programming languages.
@@ 1.4 Approximate inference for NR by variational inference
HMC has shown good performance across a range of inference scenarios. However, in the limit of very large data and many unknown parameters, even advanced sampling schemes can suffer from slow convergence and high computational complexity. In such settings, variational approaches (collectively known as variation inference (VI)) have shown significant promise by approximating the true posterior with a parametric distribution [30]. VI essentially transforms the inference problem back into an optimisation problem that can be solved by gradient descent methods.
@@
Where $\phi$ are some deterministic unknown parameters that control the shape of a surrogate distribution $q$. It is important here that the choice of $q$ is a valid distribution (that it integrates to unity and covers the support of the posterior) regardless of the parameters $\phi$. For this reason it is common to choose families of probabilities distributions or mixture models for $q$ [31, 32].
@@
Given that we want to minimise the KL divergence between $q$ and $p$ and that it is invariant to changes in $\phi$ we can essentially ignore this term in the optimisation and focus on the expectation in the above. This expectation term is often called the ‘evidence lower bound’ (ELBO) because it provides a lower bound on the unknown evidence term. This can be proved with Jensen’s inequality for probability distributions [30]. For a concave function $f$,
@@
Note that these expectations are rarely available analytically. In practice, they are approximated by Monte-Carlo. A common choice in the VI literature is to make this approximation using a single sample [33].
@@
It has been observed empirically that single sample Monte-Carlo approximations to the ELBO have low variance and give good performance [33]. However, care must be taken to ensure that the gradients remain available through the sampling procedure. A common approach is to reparametrise $q$ such that the sampling procedure does not depend on $\phi$ [^1] For more detail on the reparameterisation trick see [33].. Note that in the context of NR the above loss gradient once again requires access to the gradient of the forward reflectivity model, made possible by the contribution of this work. With access to the gradient of the ELBO all that remains is to solve the optimisation problem with a gradient descent scheme such as one of those described in section 1.2. It is important to note here that VI based on minimisation of the KL divergence can display ‘mode-seeking’ behaviour that emphasises high probability regions of the posterior at the expense of lower probability areas leading to underprediction of variance. For a discussion of why this occurs and advice in the context of NR see the appendices of this paper.
In the limit of very large data (>10k samples), computing the expectations above (even with a single sample approximation) may become prohibitively expensive. In such contexts, a particularly powerful extension is Stochastic variational inference (SVI) [31]. It is the opinion of the authors that SVI is particularly promising in the large data, complex model regime of NR having been successfully applied to datasets numbering millions of examples in other fields [31]. However, because most static NR datasets comprise <1k samples, only a standard VI approach is considered in this work.

### v2

# Towards real-time surrogate-free Bayesian inversion for neutron reflectometry

@@ Abstract
Neutron reflectometry (NR) is a key enabling technology for many areas of scientific development. Although the forward reflectivity model is well-known, inferring the physical properties of a sample from NR data requires the solution of an inverse problem. Increasingly, beamline scientists are using NR in fast kinetic configurations and probing highly-complex structures and interfaces, introducing significant uncertainty. Existing uncertainty quantification (UQ) approaches in NR, such as Markov-Chain Monte-Carlo, suffer from poor sample efficiency and slow convergence times. Recently, surrogate machine learning models have been proposed as an alternative. However, physical intuition is lost when replacing governing equations with fast surrogates. Instead, we propose a rapid, surrogate-free Bayesian inversion for NR. Our approach offers a step-change in inference speed and efficiency. For the first time in NR, exact gradients through the reflectivity are computed, enabling highly performant gradient-based inference schemes: Hamiltonian Monte-Carlo offers significant advances in sample efficiency compared to MCMC. Variational inference enables approximate UQ on the order of seconds rather than hours. We demonstrate state-of-the-art performance on a thick oxide quartz film, and robust co-fitting performance in the high complexity regime of organic LED multilayer devices. Additionally, we provide an open-source library of reflectometry kernels in the python language.
Keywords Neutron Reflectometry $\cdot$ Bayesian Inference $\cdot$ Automatic differentiation
@@ 1 Introduction
Neutron reflectometry (NR) is a powerful technique to probe surfaces and interfaces. As NR is inherently an indirect measurement technique, access to the physical quantities of interest (layer thickness, scattering length density, roughness), necessitate the solution of an inverse modelling problem, that is inefficient for large amounts of data or complex multiplayer structures (e.g. lithium batteries / electrodes). This inefficiency is particularly pronounced when uncertainty quantification is desired. Sampling schemes such as Markov-chain Monte-Carlo (MCMC) can suffer from slow convergence and poor sample efficiency as the number of unknown parameters increases. Recently, surrogate machine learning models have been proposed as an alternative to existing optimisation and inference routines. Although such approaches have been successful in some areas, physical intuition is invariably lost when replacing governing equations with fast neural networks or other surrogates.
In this paper, we seek to address the shortcomings of existing approaches to uncertainty quantification (UQ) in NR. The key innovation is that we access the gradients of the forward specular reflectivity of a thin multilayer film w.r.t the unknown parameters of interest, unlocking a host of powerful optimisation and inference techniques that remain thus-far unexploited in the context of neutron reflectometry. The principal advantages of such an approach are:
  • Order of magnitude gains in sample-efficiency (using Hamiltonian Monte-Carlo) and inference speed (using a variational approach) compared to gradient-free MCMC approaches.
  • The exact forward reflectivity laws are used to perform the inversion, no approximate pre-trained surrogate model is required.
  • Full Bayesian posterior that can accommodate arbitrary prior knowledge.
@@
The proposed approach is demonstrated in two NR case study examples: A benchmark Quartz film and a co-fit to four LED devices. Additionally, an open-source software library is provided for fast computation of forward reflectivity and gradients in the python language at github.com/MDCHAMP/refjax.
1.1 Background
NR is a widely used technique for the nanometer scale characterisation of soft matter [skoda2019recent], thin films [kirby2012phase] and bio-materials [junghans2015analysis]. Increasingly, this has involved the study of more complex layered structures and ones that are evolving in time (dynamic). The application of NR has lead to many high-profile breakthroughs in the physical and biological sciences. Polarized neutron reflectivity (PNR) was crucial for understanding the oscillatory exchange coupling behaviour so fundamental to giant magnetoresistance (GMR) [Majkrzak1988], a pivotal modern technology. Ongoing work has investigated how surfactant mixtures perform at varying temperatures using surface tension and neutron reflectivity [liley2019performance], this has been crucial for developing more efficient, low-temperature detergents, vital for reducing the carbon footprint of washing laundry. Ongoing NR studies have also been important for the study of biological membranes [pabst2010applications], their interaction with drugs [clifton2015accurate], such as antimicrobial peptides and other antibiotics.
NR is a scattering technique that illuminates a sample of interest using thermal neutrons produced and moderated from either a reactor or a spallation neutron source. NR measures the specular reflectivity $R$ as a function of the momentum transfer $Q$ from counting neutrons using a detector [penfold1990application]. Theoretically, the forward reflectivity model of complex materials is well understood. Given the physical and apparatus parameters $\theta$, one can appeal to Fresnel reflectivity to compute the predicted specular reflectivity $\hat{R}$. NR is however, inherently an indirect technique, which requires the reflectivity data to be modelled, as it is the absolute square of Fourier transform of the scattering length density, which gives rise to a loss in phase information. The parameters of interest cannot be directly accessed from measurements of the reflectivity. Fitting a parametric model of the sample is the inverse problem at the core of NR measurements and data analysis.
In the four decades since the widespread adoption of NR as a measurement technique, many methods have been proposed to solve the inverse modelling problem. Until recently, advances in NR analysis and fitting have kept close pace with the progress in machine learning (ML). The rapid rise of deep learning, automatic differentiation and specialised computation libraries has revolutionised the scalability of ML in many other fields. Yet, these analysis routes have so far been very slow to translate to NR fitting. Among the leading software packages for NR and thin-film analysis [nelson2019refnx, woollam2017completeease], the underlying fitting routines are often based on traditional data optimisation techniques such as Levenberg-Marquardt (LM) [levenberg1944method, marquardt1963algorithm], Broyden-Fletcher-Goldfarb-Shanno (BFGS) [nocedal2006numerical], or heuristic techniques such as differential evolution [storn1997differential] or particle swarms [eberhart1995new]. Other approaches follow Bayesian methodologies such as Markov-chain Monte-Carlo (MCMC) [metropolis1953equation, hastings1970monte] or nested sampling [skilling2004nested]. The authors are not currently aware of any widely-used NR software that fully utilises the power of modern automatic-differentiation and gradient-based optimisation techniques to directly solve the inverse problem.
Although no direct applications of automatic-differentiation are to be found in the literature, surrogate models have very recently been proposed to replace the inverse problem [doucet2021machine, hinderhofer2023machine, pithan2023closing, starostin2025fast] with ML. These ML models supplant the forward reflectivity model and are trained to predict the parameters of interest directly from the measured $R$. Although such approaches are able to leverage advanced ML technologies, it is inevitable that physical intuition is lost when replacing governing equations with fast machine learning.
1.2 The NR inverse problem
@@
At the very core of NR data analysis, is the inverse problem of obtaining estimates of unknown parameters of interest pertaining to both the slab model of the sample $\theta_{\text{slab}}$ (layer thicknesses, scattering length densities and roughnesses) and the instrument $\theta_{\text{NR}}$ (scaling/alignment parameters and the background detection level) (Figure 1). For notational simplicity, in the description that follows these unknown parameters shall be referred to jointly as $\theta=\{\theta_{\text{NR}},\theta_{\text{slab}}\}$. The Abeles matrix formalism [abeles1949theorie] provides a physical model of the reflectivity in terms of these parameters and the momentum transfer $\hat{R}(Q,\theta)$. Predictions from the forward model can be compared to the observed reflectivity in order to determine the values of the unknown parameters. A key challenge in NR, is the non-uniqueness of the inverse problem because the phase information is lost when measuring the reflected neutron intensity. For a particular measured neutron reflectivity measurement $R(Q)$, there may be many possible slab model solutions and therefore many sets of parameters that describe the data equally well. In practice, this is circumvented by initialising optimisation routines in the vicinity of optima that are physically plausible given prior knowledge of the sample under investigation. Such local optimisation is an area in which gradient-based methods excel, particularly in high-dimensions.
@@
In both the optimisation and inference setting, powerful techniques based on automatic differentiation exist in the ML literature for solving inverse problems. Access to the gradients of the error function (from this work, Figure 2) means that these approaches can be readily applied to NR. The remainder of this section will be dedicated to introducing approaches from both categories that are applied in the case studies outlined in this paper.
@@ 1.3 Optimisation for NR by gradient descent
Although several approaches that use approximate gradient information [levenberg1944method, marquardt1963algorithm, nocedal2006numerical] are common in the NR literature, the authors are not aware of any method that explicitly harnesses the analytical gradient with respect to the parameters of interest using automatic differentiation techniques. Methods in this class have seen intensive development since the rise of deep neural networks, particularly stochastic gradient descent methods. Access to gradient information as in this paper will make the application of these schemes very straightforward to the NR community. Efficient implementations of many popular algorithms are readily available in a number of programming languages. The most popular methods (e.g. Adam [kingma2014adam], RMSPROP [hinton2012rmsprop]) make use of gradient averages between iterations and momentum terms that help to overcome local minima. Such approaches help to alleviate some of the inherent shortcomings of these popular gradient descent approaches when applied to NR data analysis, such as becoming trapped in local minima and computational efficiency in the case of high dimensions of very many parameters.
@@ 1.4 Inference for NR by Hamiltonian Monte-Carlo
In the interest of keeping this paper self-contained and in order to appeal to practitioners of NR who may not be familiar with gradient-based sampling and approximate inference schemes, brief overviews of Hamiltonian Monte-Carlo (HMC) and variational inference (VI) are included here. The interested reader is directed to [betancourt2017conceptual] (HMC) and [blei2017variational] (VI) for more comprehensive tutorial overviews.
@@
Where the differentiation has removed the unknown normalisation constant in the kinetic energy and the ODE now depends on the hyperparameters, likelihood and prior only. As with any other nonlinear ODE it is necessary to integrate forward in time numerically in order to obtain trajectories (and therefore samples from the target posterior distribution). Symplectic integrators are a class of integrator specifically designed to simulate Hamiltonian dynamics, by attempting to conserve (or bound the drift in) the systems total energy through the Hamiltonian. A popular solver in the context of HMC is the leapfrog algorithm [hairer2006geometric]. For a given integration period $\Delta t$, the dynamics are updated sequentially as,
@@
Thus, the overall sampling scheme consists of a Markov chain with the following kernel. In each iteration, an initial momentum is drawn, and the integration is run for a number of steps $L$ with a step size of $\Delta t$. Although the position at the end of the integration window should theoretically be a valid sample from the target distribution, in practice, numerical errors in the integration can lead to bias. To overcome this, a Metropolis-Hastings step is typically taken after integration. At the end of the integration window, the last value of the parameter is accepted as the next state of the Markov Chain with probability $\alpha$. It can be shown (in [betancourt2017conceptual] for example) that the correct choice of $\alpha$ to ensure detailed balance in the Markov Chain is given by,
@@
In practice, the performance of the HMC scheme comes to depend strongly on the choice of the hyperparameters $L,\Delta t$ and $M$. However, the adaptive specification of these parameters is a well studied problem. Variants of the standard HMC algorithm such as the no U-turn sampler (NUTS) [hoffman2014no] have shown incredible robustness to a wide range of posterior geometries. NUTS is now a standard approach in Bayesian inference and implementations can be found in many programming languages.
@@ 1.5 Approximate inference for NR by variational inference
HMC has shown good performance across a range of inference scenarios. However, in the limit of very large data and many unknown parameters, even advanced sampling schemes can suffer from slow convergence and high computational complexity. In such settings, variational approaches (collectively known as variation inference (VI)) have shown significant promise by approximating the true posterior with a parametric distribution [jordan1999introduction]. VI essentially transforms the inference problem back into an optimisation problem that can be solved by gradient descent methods.
@@
Where $\phi$ are some deterministic unknown parameters that control the shape of a surrogate distribution $q$. It is important here that the choice of $q$ is a valid distribution (that it integrates to unity and covers the support of the posterior) regardless of the parameters $\phi$. For this reason it is common to choose families of probabilities distributions or mixture models for $q$ [hoffman2013stochastic, yin2018semi].
@@
Given that we want to minimise the KL divergence between $q$ and $p$ and that it is invariant to changes in $\phi$ we can essentially ignore this term in the optimisation and focus on the expectation in the above. This expectation term is often called the ‘evidence lower bound’ (ELBO) because it provides a lower bound on the unknown evidence term. This can be proved with Jensen’s inequality for probability distributions [jordan1999introduction]. For a concave function $f$,
@@
Note that these expectations are rarely available analytically. In practice, they are approximated by Monte-Carlo. A common choice in the VI literature is to make this approximation using a single sample [kingma2014auto].
@@
It has been observed empirically that single sample Monte-Carlo approximations to the ELBO have low variance and give good performance [kingma2014auto]. However, care must be taken to ensure that the gradients remain available through the sampling procedure. A common approach is to reparametrise $q$ such that the sampling procedure does not depend on $\phi$ [^1] For more detail on the reparameterisation trick see [kingma2014auto].. Note that in the context of NR the above loss gradient once again requires access to the gradient of the forward reflectivity model, made possible by the contribution of this work. With access to the gradient of the ELBO all that remains is to solve the optimisation problem with a gradient descent scheme such as one of those described in section 1.3. It is important to note here that VI based on minimisation of the KL divergence can display ‘mode-seeking’ behaviour that emphasises high probability regions of the posterior at the expense of lower probability areas leading to underprediction of variance. For a discussion of why this occurs and advice in the context of NR see the appendices of this paper.
In the limit of very large data (>10k samples), computing the expectations above (even with a single sample approximation) may become prohibitively expensive. In such contexts, a particularly powerful extension is Stochastic variational inference (SVI) [hoffman2013stochastic]. It is the opinion of the authors that SVI is particularly promising in the large data, complex model regime of NR having been successfully applied to datasets numbering millions of examples in other fields [hoffman2013stochastic]. However, because most static NR datasets comprise <1k samples, only a standard VI approach is considered in this work.